### v2

# VideoChat-Flash: Hierarchical Compression for Long-Context Video Modeling

Abstract
Long-context modeling is a critical capability for multimodal large language models (MLLMs), enabling them to process long-form contents with implicit memorization. Despite its advances, handling extremely long videos remains challenging due to the difficulty in maintaining crucial features over extended sequences. This paper introduces a Hierarchical visual token Compression (HiCo) method designed for high-fidelity representation and a practical context modeling system VideoChat-Flash tailored for multimodal long-sequence processing. HiCo capitalizes on the redundancy of visual information in long videos to compress long video context from the clip-level to the video-level, reducing the compute significantly while preserving essential details. VideoChat-Flash features a multi-stage short-to-long learning scheme, a rich dataset of real-world long videos named LongVid, and an upgraded “Needle-In-A-video-Haystack” (NIAH) for evaluating context capacities. In extensive experiments, VideoChat-Flash shows the leading performance on both mainstream long and short video benchmarks at the 2B and 7B model scale. It firstly gets 99.1% accuracy over 10,000 frames in NIAH among open-source models.
00footnotetext: * Equal contribution. $\dagger$ Corresponding authors.
1 Introduction
Long-context video modeling is one of the most essential capabilities in multimodal large language models (MLLM). With it, MLLMs can handle hours of movies or documentaries, online long video streams, and so on using million-level video token processing.
Figure 1: Comparison with mainstream MLLMs for long videos. VideoChat-Flash improves long video understanding efficiency and effectiveness by hierarchical compression and a short-to-long learning approach, respectively.
Through video-language alignment and instruction tuning, existing video MLLMs [58, 59, 29, 31, 72, 13, 78, 34, 65, 27, 4, 26] have made significant progress in minute-level video understanding, including video captioning and question answering. However, it remains highly challenging to build MLLMs for processing extremely long videos (lasting for hours or even longer). The most crucial difficulty lies in how to enable MLLMs to efficiently understand the multimodal long context brought by long videos.
Figure 2: Comparison results on various generic video-linguistic tasks
Inspired by large language models with long context, modeling multimodal long context is widely studied from several perspectives. Some work [47, 66] represented by Gemini-1.5-Pro [47] address it by training well-performed MLLMs on long-form corpus e.g. lengthy text and thousands of frames from videos, minimizing the gap between the evaluation and learning. To make the long training and inference work possibly and efficiently, building an effective training system with multimodal long context parallelism and data packing is crucial. Existing works evolve from the basic tensor parallelism to ring-style attention implementation and to current more comprehensive solutions integrating several advanced techniques. Although the progress in system construction and hardware has made it possible to train and infer with super-long multimodal contexts, such super-long multimodal contexts have significantly reduced the training and inference efficiency of models. (For Gemini-1.5-Pro [47], a one-hour video will be converted into 921,600 tokens). Meanwhile, the high redundancy in visuals makes it particularly difficult for models to understand. Some previous efforts [52, 15, 63] have been made to compress video tokens in order to achieve higher training and inference efficiency for long videos. Besides of these accomplishments, challenges still remain that increasing (or even maintaining) context understanding is often compromised by extending its length because compression on visuals trades off partial detail representations more or less, while the data and learning discrepancies between inference and training are hard to neutralize. Consequently, some current long video models even perform worse than some image-based MLLMs on video understanding benchmarks.
In this paper, we propose a Hierarchical video token Compression method (HiCo) with high-fidelity as well as a practical context modeling system VideoChat-Flash for multimodal long sequence processing. For HiCo, it widens context window using high-density token representations by compressing tokens hierarchically. We sequentially handle long videos, segmenting them into short clips, and feed clips into a MLLM by order. Similar to previous studies, HiCo condenses the visual tokens extracted from a clip to a few one based on spatiotemporal redundancies first. Linking these compressed visual tokens from clips and user query tokens together, HiCo then progressively reduce token quantity when going deeper in LLM processing, exploiting semantic correlations between clips and world knowledge embedding for compact representations.
With HiCo, we design a video MLLM VideoChat-Flash with a multi-stage short-to-long learning scheme and a corresponding data receipt with real-world long videos. In the training strategy, we start with short videos, and their captions and question-answers (QAs) for supervised fine-tuning, shift to long ones and their textual annotations, and finally employ a mixed corpus with both short and long videos for learning. We note short videos not only enhance basic visual perception, but also help MLLMs express extremely long visuals more concisely and precisely using compression, as validated in Sec 4.3. In data, we give a new long video dataset for supervised fine-tuning (SFT) in 300,000 hours and 2 billion words of textual annotations, in addition to previous ones [57, 66, 50]. Also, we extend the current “needle-in-a-haystack” (NIAH) from single-hop to multi-hop to ensure testing results correlate with models’ context modeling other than luck.
We contribute to the long video understanding as follows:
$\bullet$ We propose a new hierarchical compression scheme to encode long videos using much fewer tokens compared to existing methods. It addresses both visual and semantic redundancies, achieving a compute reduction by two orders of magnitude without compromising its long video understanding. In evaluations, our 7B model yields state-of-the-art performance with 99.1% retrieval accuracy over 10,000 frames in the typical “Needle-in-a-haystack” among open-sourced MLLMs. Also, it outperforms other 7B models in both short and long video evaluation benchmarks.
$\bullet$ We give an effective long video understanding MLLM VideoChat-Flash. It is trained with a three-stage short-to-long strategy. To support its scalable learning, we construct the corresponding long video data for supervised fine-tuning. Our training infrastructure includes a distributed system that supports high-degree sequence parallelism, ensuring efficient and robust model training.
$\bullet$ To concisely evaluate multimodal context capacities, we extend current NIAH in videos to a multi-hop configuration. It demands MLLMs to locate the target via a chain of visual hints and answer it, characterizing models’ context memorization and understanding rigorously.
2 Related Works
Figure 3: Framework of VideoChat-Flash with hierarchical token compression (HiCo).
Multimodal Large Language Models for Long Video Understanding.
Recent advancements in multimodal large language models (MLLMs) [47, 52, 63, 57, 38, 50, 66, 51] have shown significant promise in long video understanding. To address the challenge of processing long videos, researchers focus on two key strategies: (1) extending the context window of the LLM [47, 75, 66, 62] and (2) compressing the video tokens [33, 15, 63, 53, 52, 51]. For context extension, LongVILA [66] proposes a parallel processing system that distributes computation across temporal and tensor dimensions to optimize load balancing and inter-node communication. Although the approach of expanding the context window enable the possibility of long video understanding, it falls short of reducing the high computational burden and processing costs induced by long videos, thereby imposing limitations on its practical application. For token compression, Methods represented by Llama-Vid [33] use a highly compact representation while preserving key information. The high compression ratio makes it difficult for such methods to achieve excellent long video understanding performance, and they may even be inferior to some MLLMs designed for image modeling.
Long Video Benchmarking.
The evaluation of long video understanding remains a significant challenge, given the complex interplay of object recognition, temporal reasoning, memory, and higher-order cognitive abilities. Several benchmarks [46, 52, 73, 21, 71, 6, 19, 14, 32] and evaluation methodologies have emerged to address this need, primarily focusing on question-answering (QA) tasks. EgoSchema [40] leverages the Ego4D dataset to create a benchmark of 5000 multiple-choice QA questions on three-minute video clips. VideoMME [16] uses a collection of diverse online videos with human-annotated QA pairs. MLVU [81] expands the scope to include movies, TV dramas, and surveillance footage, evaluating performance on tasks such as temporal order reasoning, plot summarization, and QA. Focusing on fine-grained information retrieval, Video-NIAH [80] presents a “needle in a haystack” challenge, emphasizing the retrieval of specific details from long videos. Similarly, LongVideoBench [64] assesses both perceptual and reasoning abilities on fine-grained features within long video sequences.
3 Method
3.1 HiCo: Efficient Long Video Modeling
To handle more than thousands of frame inputs, we approach the efficient long video understanding design from the perspective of multimodal context compression in MLLM, named as hierarchical compression (HiCo). We implement HiCo in a new MLLM VideoChat-Flash. As illustrated in Fig. 3, VideoChat-Flash leverages a three-component structure: a visual encoder, a visual-language connector, and a Large Language Model (LLM). The pipeline is divided into two main stages, specifically Clip-level Compression during the encoding of long videos and Video-level Compression within the context interaction in the LLM.
Dynamic Video Sampling.
Firstly, we need to perform frame sampling on the original video. We adopt a dual sampling strategy that both helps the model develop both micro and macro-level video understanding: fine-grained motion recognition in short clips, and broader event comprehension in longer sequences. For short videos, dense frame sampling (e.g. 15 fps) is taken to capture detailed motions. For long videos, sparse frame sampling (e.g. 1 fps) is taken to focus on event understanding.
Timestamp-aware Prompt
For video MLLMs, the ability to perceive timestamps is also a crucial capability. Unlike previous works [48, 69] that rely on additional modules or designs to achieve this, we employ a simple timestamp-aware text prompt after the video context: “The video lasts for N seconds, and T frames are uniformly sampled from it.” We find that this straightforward approach is sufficient to enable the model to perceive the timestamps of the input video, achieving excellent performance on temporally sensitive tasks such as temporal grounding. (see Tab. 3 for details)
Encoding and Compression for Video Clips.
Considering the abundant redundant and repetitive information such as backgrounds and objects that exists between adjacent frames in natural videos, we sequentially encoding a long video clip by clip, and compress its visual representation through inter-frame redundancies and semantic correlations. Formally, given a frame sequence sampled from the original video, we segment it into $T$ equally sized clips. The frames of $j_{\text{th}}$ clip are transformed by a visual encoder $\mathcal{V}$ into $M$ visual tokens, denoted as $[\mathbf{v}_{i}^{j}]_{i=1,2,..,M}$. We design a video-language connector $\mathcal{F}$ composed of a compressor and an MLP projection. It compresses $M$ tokens into $N$ ones, $[\mathbf{s}_{i}^{j}]_{i=1,2,..,N}$ (w.r.t. $N<M$), before input to a LLM:
$[\mathbf{s}_{i}^{j}]_{i=1,2,..,N}=\mathcal{F}([\mathbf{v}_{i}^{j}]_{i=1,2,..,M}),$ (1)
where $\mathcal{F}$ denotes our given connector. The specific design of the visual encoder and compressor is a non-trivial decision as lots of effective candidates (e.g. Video with Q-former, Image encoders with MLP, etc.) have been explored [31, 78]. We systematically study how visual encoders and connectors are combined in different cases for the best spatiotemporal compressed representation. As shown in Fig. 4, we consider three different numbers of tokens per frame after compression (16, 49, 196) and examined image encoders, video encoders, and four popular token compression strategies: spatial downsampling [78, 12], uneven downsampling [62], spatio-temporal resampler [31, 53], and similar token merging [5, 63]. We found that video encoder, by modeling inter-frame interactions, better exploit temporal redundancy compared to image encoder. Consequently, combining a video encoder with a parameter-free similar token merge operation facilitates learning effective spatiotemporal compressed representations.
Figure 4: Comparison of different compressors. The specific comparisons and implementation details can be found in the Sec. 4.3 and the Appendix.
Progressive Visual Dropout in LLM.
Although clip-level compression has been carried out before, due to the possibility of longer-range visual redundancies in long videos (e.g. surveillance videos), and when an LLM responds to specific instructions regarding the visual input, it may not be necessary to continuously focus on the entire long video context. We consider conducting further video-level compression during the LLM inference stage. Recent works [9, 8] have explored acceleration strategies for MLLMs when processing short visual contexts. Most of them drop visual tokens based on the correlation between text tokens and visual tokens. In contrast, we have found that when an LLM processes a long video context, it pays attention to the entire long video context at the shallow layers of the LLM, while focusing on the details of certain local moments at the deep layers (see the Appendix for specific visualizations). Based on this observation, we have designed a progressive visual dropout strategy, which is divided into two stages. At the shallow layers of the LLM, we uniformly drop a small number of video tokens (i.e. uniform drop), reducing the computation while maintaining the original spatio-temporal structure of the video context. At the deep layers of the LLM, we rely on the correlation between text tokens and video tokens to retain the most critical relevant information (i.e. text-guided select). With the $NT$ visual tokens $[\mathbf{s}_{i}]^{j}$ from the long video, we further condense them in a progressive manner during the computation in LLM by dropping visual ones. Formally, we denote $[\mathbf{t}_{i}^{(0)}]_{i=1,2,..,NT}=[\mathbf{s}_{1}^{1},\mathbf{s}_{2}^{1},...,\mathbf{s}_{1}^{T},\mathbf{s}_{2}^{T},...,\mathbf{s}_{N}^{T}]$ as the input visual tokens of the LLM, and $[\mathbf{t}_{i}^{(l)}]$ as the $l_{\text{th}}$ output of these visual tokens via LLM. Regarding this,
$[\mathbf{t}_{i}^{(l)}]=[\mathbf{r}_{i}^{(l)}*\mathbf{t}_{i}^{(l-1)}]\quad\text{w.r.t.}\quad\mathbf{r}_{i}^{(l)}\in\{0,1\}$ (2)
We have found that this operation not only effectively improves the computational efficiency of the model but also enhances the understanding performance of the model by reducing irrelevant visual noise.
3.2 Multi-stage Short-to-Long Learning
We focus on using videos and their annotations for training. Unlike previous works [75, 66] that often use long-form text to extend the context window of LLMs, we argue that direct training on long-form videos minimizes the gap between training and testing, leading to better performance in downstream long video evaluations. Using a short-to-long scheme, our proposed VideoChat-Flash is trained on a mixed dataset of both short and long videos. The training data are detailed in the Appendix.
Stage-1: Video-Language Alignment.
In this stage, we freeze the visual encoder and the large language model while training the compressor and the MLP to align the language with the compressed visual features. We use 0.5 million image-text pairs and 0.5 million short video-text pairs, and sample 4 frames from each video in training.
Stage-2: Short Video Pre-training.
To enhance the model’s understanding of visual concepts, we conduct visual pre-training using 3.5 million images and 2.5 million short video-text pairs. Note most captions are refined using AI models to ensure richer and more detailed descriptions. During this stage, we sample 8 frames from each video.
Stage-3: Joint Short & Long Video Instruction Tuning.
To enable the model to handle a wide variety of video tasks, we collect 3.5 million instruction fine-tuning samples, including 1.1M images, 1.7M short videos (under 60 seconds), and 0.7M long videos (60$\sim$3600 seconds). We mix the short and long video data to ensure the model retains fine-grained understanding while expanding its comprehension of long videos. The sampling method used is the dynamic video sampling described in Section 3.1, with the number of sampled video frames ranging from 64 to 512.
Stage-4: Efficient High-Resolution Post-finetuning.
To enable the model to perceive higher resolutions, we employ a highly efficient post-finetuning strategy to adapt the original low-resolution video encoder to higher-resolution inputs. Specifically, we increase the input resolution of the video encoder from 224 to 448, freeze the LLM, and directly utilize 25% of the stage-3 data for post-finetuning the video encoding. We find that this simple, full-data strategy effectively enhances the video encoder’s adaptability to higher-resolution video inputs.
3.3 Large-scale Corpus for Long Video Training
One of the challenges in long video model training is the shortage of large-scale, high-quality data. Though recent advances have mitigated this by long-form datasets of video-text pairs, these lack the instruction-following paradigm, such as (video, instruction, answer) triplets, crucial for multimodal reasoning. To address this, we introduce a large-scale long video instruction-tuning dataset named LongVid. It comprises 114,228 long videos and 3,444,849 question-answering (QA) pairs across five different task types, supporting models to handle diverse long video scenarios.
Our proposed dataset leverages the rich diversity of existing datasets, including Ego4D [18], HowTo100M [41], HD-Vila [67], and MiraData [23], encompassing a wide range of video types: movies, egocentric videos, news, interviews, and how-to videos, and other in-the-wild videos of long duration. For data curation, we generate dense event labels for each long video. Specifically, we utilize existing high-quality short video captions (Panda-70M [10] for HD-VILA [67], CosMo [54] for HowTo100M [41], Ego4D-HCap [22] for Ego4D [18], and the original high-quality captions provided in MiraData [23]) and filter the consecutive segments that can be regroup into a long video sequence, then we construct a sequence of event labels with their corresponding timestamps for every long video based on their captions. In this process, for datasets with high-quality event-level annotations (HT-Step [1] for HowTo100M [41], Ego4D-HCap [22] for Ego4D [18]), we directly utilize them as the event labels, while for others, we extract the major event from the caption using an LLM. Finally, we construct several types of long video QA pairs based on the video captions, event labels, and the timestamps of short video segments.
The annotations of these videos are categorized into five tasks: video captioning, temporal grounding, event relation recognition, scene relation recognition, and video event counting. Among them, the temporal grounding task is to answer the accurate timestamp for one specific event, and event relation recognition is constructed by selecting the right event order in a long video. Scene relation recognition is formulated analogously to the event relationship one. It is built for videos that have background captions (like MiraData), where we can extract the major scenery labels.
3.4 Multi-hop Needle In A Video Haystack
Figure 5: An example of “multi-hop needle in a video haystack”. Frames 1 to 3 and needle are the inserted indicating images while X1, X2, and X3 and decoy pictures for distraction. MLLMs are asked to both find the needle (by choosing its right description) and then answer its related question.
Needle in a video haystack requires the model to locate an indicated image in a long video, find a target word, or answer a question based on that image. Most existing benchmarks [75, 80] typically inject a target image into a long video, making it difficult to assess genuine long-context understanding. Models might succeed by simply finding the most visually distinct frame, instead of targeted searching according to instructions. This design results in quite high accuracy (often exceeding 99%) by almost every model on these benchmarks.
Regarding this, we propose a new benchmark, termed multi-hop needle in a video haystack. It rigorously evaluates the model’s ability to process extremely long contexts by demanding it to locate a sequence of interconnected indicating images embedded within a long video, as illustrated in Fig. 5. The first indicating image could be located via the given instruction, while subsequent images are found using clues embedded within the preceding images, creating a chain. Only after successfully navigating this whole chain and identifying the final “needle” image can the model answer a question related to that image.
To avoid trivial solutions, we give two constraints to ensure the answer can only be obtained through the designed order. First, we incorporate decoy images (three in our experiments) with misleading clues, forcing the model to follow the correct chain. Second, the model must both find the needle (by choosing the right option) and answer the corresponding question correctly, minimizing the probability of success by guessing. In practice, all images are sourced from MS-COCO [36], leveraging its human-annotated captions and question-answer pairs. This multi-hop approach provides a far more robust evaluation of long context understanding in multimodal large language models (MLLMs) than the current single-hop.
4 Experiments
Implementation details.
In dynamic video sampling, we randomly sample 64-512 frames. For VideoChat-Flash, we employ UMT-L@224 [30], token merging with MLP, and Qwen2-7B as visual encoder, connector, and LLM, respectively. When processing a long video, we divide it into shorter clips, each consisting of 4 frames. Each clip is then compressed into 64 tokens, meaning that, on average, each frame is represented by 16 tokens. Regarding video-level compression, while it presents some challenges in compatibility with training acceleration strategies such as sequence parallelism (although our experiments confirm that enabling it during training does not compromise performance), we employ different configurations during inference depending on the use case. For efficiency-focused scenarios, we utilize higher drop rates and more aggressive drop methods, whereas for performance-oriented scenarios, we apply dropout only in the later layers of the LLM.
Model FLOPs Memory(G)
(T) Train Infer
Input 64 frames
LongVILA [66] 455.1 56.7 18.1
LongVA [75] 316.8 50.6 17.4
VideoChat-Flash (ours) 29.6 35.2 15.5
Input 256 frames
LongVILA [66] 2946.3 oom 26.8
LongVA [75] 1870.1 oom 23.7
VideoChat-Flash (ours) 113.3 40.2 16.1
Input 1000 frames
LongVILA [66] 28707.7 oom 79.7
LongVA [75] 16583.7 oom 48.3
VideoChat-Flash (ours) 519.7 59.4 18.5
Input 10000 frames
LongVILA [66] 2368810.0 oom oom
LongVA [75] 1289490.0 oom oom
VideoChat-Flash (ours) 15786.0 oom 47.4
Table 1: Comparison of FLOPs and Cuda memory. FLOPs and memory are estimated by the context length of the input LLM.
4.1 General Video Understanding Evaluation
Model Size #tokens MVBench PerceptionTest LongVideoBench MLVU VideoMME (w/o & w sub.) LVBench per frame Avg Val Val M-Avg Overall Long Avg Avg. Duration 16s 23s 473s 651s 1010s 2386s 4101s Proprietary Models GPT4-V [43] - - 43.7 - 59.1 49.2 59.9/63.3 53.5/56.9 - GPT4-o [44] - - 64.6 - 66.7 64.6 71.9/77.2 65.3/72.1 30.8 Gemini-1.5-Pro [47] - - 60.5 - 64.0 - 75.0/81.3 67.4/77.4 33.1 Small Size MLLMs Qwen2-VL [55] 2B - 63.2 - - - 55.6/60.4 - - InternVL2.5 [11] 2B 256 68.8 - 46.0 61.4 51.9/54.1 - - VideoChat-Flash @448 2B 16 70.0 70.5 58.3 65.7 57.0/63.9 44.9/54.0 42.9 Open-Source MLLMs VideoLLaMA2 [13] 7B 72 54.6 51.4 - 48.5 47.9/50.3 42.1/43.8 - VideoLLaMA2 [13] 72B 72 62.0 57.5 - - 62.4/64.7 57.6/59.0 - VideoChat2-HD [31] 7B 72 62.3 - - 47.9 45.3/55.7 39.8/53.9 - InternVideo2-HD [59] 7B 72 67.2 63.4 - - 49.4/ - - - IXComposer-2.5 [74] 7B 400 69.1 34.4 - 37.3 55.8/58.8 - - InternVL2 [12] 8B 256 65.8 - 54.6 64.0 54.0/56.9 - - InternVL2 [12] 76B 256 69.6 - 61.1 69.9 61.2/62.8 - - InternVL2.5 [11] 8B 256 72.0 - 60.0 68.9 64.2/66.9 - - Qwen2-VL [55] 7B - 67.0 62.3 - - 63.3/69.0 - - Qwen2-VL [55] 72B - 73.6 68.0 - - 71.2/77.8 - 41.8 LLaVA-NeXT-Video [78] 7B 144 53.1 48.8 49.1 - - /46.5 - - LLaVA-OneVision [27] 7B 196 56.7 57.1 56.3 64.7 58.2/61.5 - - LLaVA-OneVision [27] 72B 196 59.4 66.9 61.3 68.0 66.2/69.5 - - LLaVA-Video [79] 7B 676 58.6 67.9 58.2 70.8 63.3/69.7 - - Open-Source Long Video MLLMs LLaMA-VID [33] 7B 2 41.9 44.6 - 33.2 25.9/ - - 23.9 Kangaroo [38] 8B 256 61.0 - 54.8 61.0 56.0 / 57.6 46.7 / 49.3 39.4 LongVILA [66] 7B 196 67.1 58.1 57.1 - 60.1/65.6 47.0/52.1 - LongVA [75] 7B 144 - - - 56.3 52.6/54.3 46.2/47.6 - LongLLaVA [57] 9B 144 49.1 - - - 43.7/ - 36.4/ - - LongVU [50] 7B 64 66.9 - - 65.4 - /60.6 - /59.5 - VideoChat-Flash @224 7B 16 73.2 75.6 64.2 74.5 64.0/69.4 53.6/61.9 47.2 VideoChat-Flash @448 7B 16 74.0 76.2 64.7 74.6 65.3/69.7 55.4/63.3 48.2
Table 2: Results on comprehensive video question answer benchmarks
Benchmark.
We evaluate our model on five general video understanding benchmarks in question-answering format, including two short video benchmarks: MVBench [31] and Perception Test [45], and three long video benchmarks: LongVideoBench [64], MLVU [81] and LVBench [56], and a comprehensive benchmark, VideoMME [16], covering videos ranging from minute-level to hour-level durations. We further evaluate the model’s timestamp awareness using the Charades-STA temporal grounding dataset [17].
Short video understanding.
As shown in Tab. 2, our VideoChat-Flash gives the best results on MVBench (74.0%) and Perception Test (76.2%) among all, surpassing other approaches by a notable margin. We have achieved new state-of-the-art (SOTA) results among 7B models. Moreover, our performance even surpasses that of models with much larger scales, such as InternVL2-76B and VideoLLaMA2-72B, and proprietary models like GPT4o and Gemini-1.5-Pro. Furthermore, our model achieved a comprehensive lead in results at the 2B parameter scale, even surpassing some models with 7B or even 72B parameters. This demonstrates the effectiveness of our approach in enhancing video perception capabilities.
Long video understanding.
Tab. 2 gives that VideoChat-Flash not only outperforms previous open-source general MLLMs of the same model scale but also surpasses a host of MLLMs designed specifically for long videos on the long video understanding benchmarks. Especially, it gets the new SOTA performance (74.7%) on MLVU, even compared with GPT4-V and GPT4-o. This empirically proves that VideoChat-Flash with HiCo can handle long multimodal contexts effectively. More importantly, despite performance, the training and inference efficiency of our model is also much higher than that of previous ones. As in Tab. 1, even when processing short videos, the compute load of our model is only one-tenth that of previous models. Meanwhile, as the number of input frames increases, the difference becomes more and more pronounced. On a single A100 device, only our model can complete the inference on 10,000 frames (approximately equivalent to a 3-hour video). Concretely, VideoChat-Flash’s compute load is two orders of magnitude lower than that of LongVILA [66] (15,786 vs. 2,368,810 TFLOPs).
Method Charades-STA
$R@1_{(IOU=0.3)}$ $R@1_{(IOU=0.5)}$ $R@1_{(IOU=0.7)}$
Specialized Models
2D-TAN [77] 57.3 45.8 27.9
MMN [61] 65.4 53.3 31.5
MDETR [25] 65.8 52.1 30.6
UniVTG [35] 72.6 60.2 38.6
Open-Source MLLMs (7B size)
VTimeLLM [20] 51.0 27.5 11.4
HawkEye [60] 50.6 31.4 14.5
TimeChat [48] - 32.2 13.4
TimeSuite [69] 69.9 48.7 24.0
VideoChat2 [31] 9.6 3.4 1.4
LLaVA-OneVison [27] 31.2 13.5 5.2
VideoChat-Flash@224 74.5 53.1 27.6
VideoChat-Flash@448 72.5 51.4 26.4
Table 3: Results on temporal grounding task. Similar to previous Video MLLM approaches [48, 69], our model is trained without any grounding data from the Charades-STA dataset, evaluating its zero-shot capabilities.
Temporal Grounding
As detailed in Tab. 3, in contrast to previous approaches [48, 69] that rely on intricate architectures for timestamps awareness, our model exhibits surprisingly strong temporal grounding capabilities. Its zero-shot performance substantially exceeds that of existing MLLMs, even reaching or surpassing the performance of dedicated, fully supervised models.
4.2 Long Context Evaluation with NIAH
Baseline.
LongVA [75] and LLama-VID [33] are used as baselines. LongVA trains MLLMs using long text data, transfering the long context of LLM from text to video. LLama-VID accomplishes efficient inference of long videos by compressing each frame to only two tokens. Our model benefits from these two, so we take them as baselines.
Single-Hop NIAH.
Fig. 6 presents the results of baselines and ours on typical “single-hop Needle-in-a-haystack” for long videos, widely used in long video evaluations [75, 66, 57, 50]. Adhering to the protocols in LongVA [75], we source a long video and sample frames uniformly from it. Then we add needles (indicating images) into the sampled image sequence at different positions. MLLMs are fed with this long image sequence and answer the corresponding questions to the indicating images. Different from existing video NIAH with thousands of frames, we evaluate all models over 10,000 frames. Note our VideoChat-Flash delivers a 99.1% success rate in accurately retrieving the correct indicating image and answering the related question even across 10,000 frames. In comparison, LongVA gives a decent result close to 92% within 3000 frames while LLama-VID only achieves 55% accuracy. It demonstrates VideoChat-Flash’s state-of-the-art performance in long multimodal context modeling.
(a) LongVA [75], accuracy=91.8% at 3k frames
(b) LLama-VID [33], accuracy=55.0% at 10k frames
(c) VideoChat-Flash (ours), accuracy=99.1% at 10k frames
Figure 6: Results on the single-hop Needle-in-a-haystack evaluation with 10,000 frames.
Figure 7: Results on the multi-hop Needle-in-a-haystack with 10,000 frames.
Multi-Hop NIAH.
Fig. 7 shows that VideoChat-Flash beats baselines non-trivially in multi-hop NIAH. In evaluations, MLLMs need to trace along the chain of indicating images, locate the needle, and answer its question. Two metrics “CAP” and “’QA’ are used to denote the accuracy of finding the correct needle and the accuracy of answering the questions related to the needle as well as finding the needle successfully, respectively. As in the single-hop setting, we also sample 10,000 frames for testing. Specifically, VideoChat-Flash gives 31.3% and 25.4% in “CAP” and “QA” on average, higher than LongVA by around 8 points. Note that LongVA only works within 3000 frames. Compared with single-hop experiments, this shows VideoChat-Flash’s high-performing multimodal context in memorizing and understanding the quite long frame sequence.
4.3 Ablation & Analysis
We ablate key designs including hierarchical compression in VideoChat-Flash as follows.
Dynamic video sampling strategy.
As shown in Tab. 4, setting a lower bound of sampling frame number during training can significantly improve the fine-grained understanding in short videos. When the frame number increases from 8 to 64, the performance on MVBench is improved by 2.3%. Moreover, a varying sampling number for long videos is superior to a fixed number.
Training Testing MVBench VideoMME
Min Max Min Max Avg Overall
8 64 64 64 68.1 58.0
32 64 64 64 70.2 57.9
64 64 64 64 71.0 58.3
64 128 64 64 71.4 58.8
64 128 64 128 71.7 59.3
64 128 128 128 70.9 60.3
128 128 64 64 70.9 58.0
128 128 128 128 71.4 58.3
Table 4: Ablation of video sampling method.
Video encoders are efficient clip encoders.
Recent video MLLMs almost all use image encoders (e.g. SigLip) for video encoding. How image encoders and video ones perform in the context of video understanding is relatively under-explored. We empirically compare them in both short and long video evaluations and confirm that video encoder is still effective in encoding clips. As given in Tab. 5, we test the most popular image encoder, SigLIP [70], and the short video encoder, UMT [30], under different compressors. We find that even when the computational load is only a quarter of that of SigLIP, UMT still demonstrates remarkable performance and performs better on MVBench, demonstrating the merit of a video encoder in detailed temporal modeling.
Model FLOPs Latency MVBench VideoMME
(G) (ms) Avg Overall
w/ Spatial Downsampling, #tokens per frame=49
SigLIPSO400M@384 [70] 2679 79.7 60.2 53.6
UMT-L@224 [30] 596 11.8 61.5 53.1
w/ Resampler, #tokens per frame=49
SigLIPSO400M@384 [70] 2679 79.7 55.5 51.5
UMT-L@224 [30] 596 11.8 59.5 51.9
w/ Token Merging, #tokens per frame=49
SigLIPSO400M@384 [70] 2679 79.7 61.4 55.3
UMT-L@224 [30] 596 11.8 61.9 53.6
w/ Token Merging, #tokens per frame=16
SigLIPSO400M@384 [70] 2679 79.7 60.2 53.5
UMT-L@224 [30] 596 11.8 61.1 53.2
Table 5: Comparison of visual encoders.
Comparison of different visual compressors.
As shown in the Tab. 6, we found that the simple spatial downsampling performs better than the spatiotemporal resampler under various compression ratios. This indicates that it might be rather difficult to obtain a good compressor in a learnable manner with the currently relatively low amount of training data. The uneven downsampling, which utilizes the prior knowledge of inter-frame redundancy, can further enhance the effect of spatial downsampling. Meanwhile, directly merging similar tokens by leveraging the similarity among tokens is the most effective. When combined with the unfrozen visual encoder during training, a good compact spatiotemporal representation can be learned, and it can even outperform the results without compression when the compression ratio is low.
Compreesor MVBench MLVU VideoMME Avg
Avg M-Avg Overall
#tokens per frame=729
Baseline (Uncompressed) 59.4 64 55.3 59.6
#tokens per frame=196
Spatial Downsampling 60.2 63.7 52.8 58.9
Uneven Downsampling 60.9 62.5 54.9 59.4
Spatio-temporal Resampler 59.5 61.9 51.9 57.8
Similar Token Merging 62.8 66.7 56.8 62.1
#tokens per frame=49
Spatial Downsampling 60.2 61.8 53.6 58.5
Uneven Downsampling 59.8 62.8 54.3 59.0
Spatio-temporal Resampler 55.5 58.1 51.1 54.9
Similar Token Merging 61.4 63.3 55.3 60.0
#tokens per frame=16
Spatial Downsampling 58.1 61.1 50.1 56.4
Uneven Downsampling 58.3 60.0 52.3 56.9
Spatio-temporal Resampler 51.4 54.7 47.7 51.3
Similar Token Merging 60.2 62.4 53.5 58.7
Table 6: Comparison of compressors.
Progressive visual dropout.
As shown in the Tab. 7, at the shallow layers of the LLM, uniform dropout performs better than attention select on long video tasks. However, at the deep layers of the LLM, attention select shows better performance. Performing visual dropout at the deep layers can not only improve the computational efficiency but also enhance the performance. Combining uniform dropout and attention select can achieve a good balance between performance and efficiency. More relevant analyses and comparative experiments can be found in the Appendix.
Drop type/keep ratio Drop layer FLOPs Latency MLVU VideoMME
num. layers=28 (G) (s) M-Avg Overall
- - 341.4 2.6 71.8 61.2
uniform/0.5 4 242.8 1.9 71.2 60.4
attention/0.5 4 242.8 1.9 70.7 60.8
uniform/0.5 18 295.2(-14%) 2.2 71.7 61.8
attention/0.5 18 295.2 2.2 72.1(+0.3) 61.7(+0.5)
attention/0.75,attention/0.25 4,18 245.8 1.9 71.4 60.9
uniform/0.75,attention/0.25 4,18 245.8(-28%) 1.9 72.0(+0.2) 61.1(-0.1)
Table 7: Effectiveness of visual dropout.
Multi-stage Short-to-Long Learning.
In Tab. 8, using a large amount of short footage for pre-training significantly improves the understanding of short videos as well as long ones. Incorporating both long videos and short videos in supervised fine-tuning (sft) of the model further enhances long video understanding while maintaining the ability to understand short videos. Additionally, adding the proposed long video data in Sec. 3.3 notably improves the performance in long video benchmark MLVU, at nearly no cost of performance drop in short video benchmark MVBench.
Settings MVBench MLVU VideoMME
Avg M-Avg Overall
Baseline (short video SFT) 61.5 61.6 53.1
$+$short video pretraining 65.8 63.1 54.3
$+$long video post-FT 67.0 64.5 55.5
$+$joint short & long video FT 67.0 64.5 55.5
$+$our long video SFT data 66.5 68.3 55.8
Table 8: Ablation of training strategies and data.
5 Conclusions
In this paper, we address the challenge of long-context video modeling in MLLMs by introducing the Hierarchical video token Compression (HiCo). HiCo effectively reduces the compute through a hierarchical strategy capitalizing redundancies in visual details and semantics, preserving vast amounts of visual essentials in a compact form. Based on HiCo, we design a practical context modeling system VideoChat-Flash. With its multi-stage short-to-long learning scheme, rich dataset of real-world long videos, and robust distributed computing framework, VideoChat-Flash ensures that MLLMs can be trained and deployed at scale for handling hour-scale videos with millions of tokens, achieving leading performance in long and short video benchmarks.
Supplementary Material
6 Implementation Details and More Analysiss
6.1 Clip Compressor
As shown in Fig. 4, we consider four popular token compression strategies to compress the features from video clips:
  • Spatial Downsampling. Applying spatial operations (pooling [65], interpolation [78], and convolution (pixel shuffle) [12]) to each video frame for downsampling has been demonstrated in previous work [65, 39] as an effective method to reduce the number of video tokens. However, due to the lack of temporal interaction, this approach fails to leverage the relation between frames. We use pixel shuffle in our experiments.
  • Uneven Downsampling. Considering the similarities between adjacent frames, it is unnecessary to retain full details for every frame. We can apply down-sampling operations with different sizes across frames within a clip. Specifically, a lower down-sampling size is applied to the first frame, while higher down-sampling sizes are used for the remaining frames. Similar approaches have been validated in a recent study [62].
  • Spatio-Temporal Resampler. Using a learnable compressor, such as a Q-Former [31] or a cross-attention layer, to compress spatiotemporal tokens. However, this approach requires a large amount of data for effective learning. In training, we observe that the Q-Former barely converges well in our setting. So in our ablations, we adopt a single-layer cross-attention instead.
  • Similar Token Merging. We directly merge similar tokens, using the ToMe [5] approach.
6.2 Visual Dropout in LLM
Figure 8: Visual redundancy in long context across layers. We conduct experiments on Qwen2-7B (28 layers) and test the impact of droping 50% of the visual tokens from shallow to deep layers.
Visual token redundancy in LLM inference.
As shown in Fig. 8, we find that even when half of the tokens are discarded at the shallow layers of the LLM, the performance of long video understanding only degrades marginally. This indicates that despite high compression at the clip level (encoding each frame into only 16 tokens), there remains considerable redundancies between clips when their representations are interacted in the LLM. Furthermore, we find the overall understanding performance gets better as the dropout happens in the deeper layer of the model. Remarkably, at approximately two-thirds of the LLM’s depth, the performance even surpasses that of the no-discard baseline. This might suggest that in the deeper layers of the network, an excess of visual tokens may interfere with the model’s reasoning process. For the drop type, we observe that uniform drop often outperforms attention-based selection in the shallow layers. We suppose, at these layers, the LLM has not yet fully determined the specific locations to focus on. As a result, relying on attention may introduce bias.
Visualization of visual attention map.
As shown in the  Fig. 9, for long video context, the attention of text tokens is relatively dispersed in the shallow layers of the network. However, as the layers deepen, the attention gradually becomes focused on specific regions. Thus, we believe that the attention scores in the deeper layers are more reliable, while those in the shallow layers may be prone to bias.
Figure 9: Visualization of the attention scores from the last textual token to visual tokens at each layer of the network.
7 Training Details
7.1 Training hyperparameters.
As shown in Table 1, the training details and hyperparameters for each stage of our VideoChat-Flash model are presented.
Stage-1 Stage-2 Stage-3 Stage-4
Vision Resolution×$\times$×Num. frames 224 224 $\times$8 224$\times$(64$\sim$512) 224$\times$(64$\sim$512)
#Tokens 16$\times$4 16$\times$8 16$\times$(64$\sim$512) 16$\times$(64$\sim$512)
Data Dataset Image & Short Video Image & Short Video (Multi)-Image & Short/Long Video (Multi)-Image & Short/Long Video
#Samples 1M 4M 3.2M 0.3M
Model Trainable Projector Full Model Full Model ViT&Projector
7.6B LLM 20.0M 7.9B 7.9B 0.3B
Training Batch Size 512 256 256 256
LR of vision encoder 1×$\times 10^{-3}$ 2×$\times 10^{-6}$ 2×$\times 10^{-6}$ 2×$\times 10^{-6}$
LR of connector & LLM 1×$\times 10^{-3}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$
Epoch 1 1 1 1
Table 9: Training details of each training stage for the VideoChat-Flash-7B model
7.2 Training Data
Stage 1: Video-Language Alignment.
In this stage, we use 558k image-text pairs from LCS-558K [37] and 481k short video-text pairs from S-MiT [42].
Stage 2: Short Video Pre-training.
To enhance the model’s understanding of visual concepts, we conduct visual pre-training using 3.5 million images and 2.5 million short video-text pairs.
  • Video Description Data. We utilize the video description data recaptioned with VideoChat2 [31] from WebVid2M [3].
  • Detailed Video Description Data. We employ the 323k detailed video description data recaptioned with Gemini [47] from WebVid [3] and Kinetics [24], as in previous work [49].
  • Detailed Image Description Data. We use the 3.5 million detailed image description data recaptioned with LLava-NeXT-34B [78] from the following datasets: COCO118K, BLIP558K, and CC3M, as provided by previous work [27].
  • Text Data. To enhance the model’s language understanding capabilities, we incorporate 143K samples from the Evo-Instruct dataset [7].
Stage 3: Joint Short & Long Video Instruction tuning.
To enable the model to handle a wide variety of video tasks, we collect 3.5 million instruction fine-tuning samples, including 1.1M images, 1.7M short videos (under 60 seconds), and 0.7M long videos (60$\sim$3600 seconds).
  • Image Instruction data. We primarily utilized single-image instruction data from LLava-NeXT [78], Allava [7], and ShareGPT4-o [12, 59]. Additionally, we incorporated multi-image data provided by LLaVA-Interleave [28].
  • Short Video Instruction data. We primarily utilized short video data from VideoChat2 [31] and InternVideo2 [59] for instruction fine-tuning. Additionally, we incorporated data annotated with GPT4-o from previous works, including ShareGPT4o [12, 59], VideoChatGPT-Plus [39], LLaVA-Video-178K [79] and LLava-Hound [76].
  • Long Video Instruction data. We primarily utilized long video instruction data from MoiveChat [52], Vript [68] and our LongVid.
8 Dataset Details of LongVid
The videos of LongVid are curated from 4 open-source video datasets: Ego4D [18], HowTo100M [41], HD-VILA [67], and MiraData [23]. We provide details of the data construction pipeline for each dataset as follows.
8.1 Ego4D
For ego-centric videos, we adopt 3,662 long videos from the Ego4d [18] and leverage Ego4DHcap [22] as the corresponding captions. Ego4DHcap gives hierarchical captions for short, medium, and long video segments. For the short video captioning task, we directly utilize these captions, while for the dense caption task, we concatenate captions in the lower level to form a dense one. For example, we merge all short video captions in a medium video segment to create a dense medium-level one, and the dense caption of long video segments can be formed by concatenating multiple medium-level video captions.
We also build event relation recognition and temporal grounding tasks based on captions of short video segments. For the event relation recognition task, models are required to choose the right order of an event sequence. Since we find the captions of short videos are highly concise and event-oriented, we use them as the event labels and serially put the short captions in a medium-level video segment as the ground-truth event relationship. For the temporal grounding task, we use the short video captions with the corresponding timestamps as the ground-truth, and randomly select other timestamps in the current medium video segments as the false options.
8.2 MiraData
MiraData [23] provides multi-level captions for large-scale minute-level movie segments. Apart from short and dense captions that are used for short and dense video captioning tasks, it also provides multiple fine-grained captions that focus on various specific perspectives, such as the main object, background, camera movements, and video style. We use an open-source LLM (Qwen-72b [2]) to extract the event and background labels from the main object and background captions, respectively, and we put the labels of a long video in the right order as the ground truth of the event/background relation recognition task. For the temporal grounding task, we use the event label with the corresponding timestamp as the ground-truth option.
8.3 HowTo100M
HowTo100M [41] includes more than 1 million long-duration how-to videos. We adopt HowToInterlink7M [54], a video captioning dataset that provides refined interleaved video captions of HowTo100M videos as short and dense video captions. For the event relationship recognition and temporal grounding tasks, we use HTStep [1], a large-scale dataset containing temporal annotations of instructional steps in HowTo100M videos.
8.4 HD-VILA
While previous datasets focus on long videos in specific domains, we also select part of the videos from HD-VILA [67], a large-scale video dataset that includes various in-the-wild videos. We argue that adding these videos into training could enhance the model’s ability to process long videos in some uncommon domains. For HD-VILA videos, we adopt the captions of Panda-70M [10]. Specifically, we filter consecutive video segments that can be re-constructed into more than 60s long videos from the 10M training subset and utilize these captions as the video short/dense captioning and temporal grounding tasks. The event labels are also extracted from these captions in the same way as MiraData [23].